[figure: 9.7 cm]

ELECTRONIC TRANSPORT THROUGH MAGNETIC NANOSTRUCTURES

A. Brataaas^{a}, R. van Gorkom^{a}, G. Tatara^{b}, P.J. Kelly^{c}, Yu.V. Nazarov^{a},

and G.E.W. Bauer^{a} aDepartment of Applied Physics and DIMES Delft University of Technology, 2628 CJ Delft, The Netherlands

bGraduate School of Science, Osaka Univsersity, Toyonaka, Osaka 560-0043, JAPAN

cFaculty of Applied Physics, University of Twente, P.O. Box 217 7500 AE Enschede, The Netherlands

Theoretical work of the authors on transport through magnetic nanostructures is reviewed. Topics to be discussed are the first-principles theory of transport through ballistic and diffuse systems, phenomenology of the Coulomb blockade in magnetic singleelectron transistors and transport through structures with non-collinear magnetization, like magnetic domain walls and spin-injection devices.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : PLENARY LECTURE

Corresponding author : Gerrit E.W. Bauer

Address for correspondence : Department of Applied Physics and DIMES Delft University of Technology, 2628 CJ Delft, The Netherlands

Email address : G.E.W.Bauer@TN.TuDelft.NL